2.1) McIver’s Meals, Inc. currently pays a R2 annual dividend. Investors believe that dividends will grow at 12% for the next year, and will have a constant growth rate of 7% after year 2. Assume the required return is 10%. What is the current market price of the stock?                                                                                       (6)

             2.2) A firm that manufactures ventilators is contemplating the replacement of its old printing   

                   machine with a new model costing R800,000. The old machine, which originally cost

                   R300,000, has 3 years of expected life remaining and a current book value of R290,000    

                   versus a current market value of R340,000. Target's corporate tax rate is 30 percent. If

                   the firm sells the old machine at market value, what is the initial after-tax outlay for the

                   new machine?                                                                                                            

           2.1) McIver’s Meals, Inc. currently pays a R2 annual dividend. Investors believe that dividends will grow at 12% for the next year, and will have a constant growth rate of 7% after year 2. Assume the required return is 10%. What is the current market price of the stock?                                                                                       (6)

             2.2) A firm that manufactures ventilators is contemplating the replacement of its old printing   

                   machine with a new model costing R800,000. The old machine, which originally cost

                   R300,000, has 3 years of expected life remaining and a current book value of R290,000    

                   versus a current market value of R340,000. Target's corporate tax rate is 30 percent. If

                   the firm sells the old machine at market value, what is the initial after-tax outlay for the

                   new machine?                                                                                                            
Answer #1

Answer : 2.1) Calculation of Current price :

Current Market Price = Present Value of Dividends and Stock Price

Below is the table showing calculation Current Market Price

Year Dividend /Price PVF @10% Present Value of Dividend
0 2 1
1 2.24 0.909090909 2.036363636
1(Price) 79.89333333 (Working Note) 0.909090909 72.63030303
Total 74.66666667

Working Note :

Price in Year 1 = Dividend in year 2 / (Required Return - Growth rate)

= [2.24 * (1.07)] / (0.10 - 0.07)

= 2.3968 / 0.03

= 79.893333

Current Price is 74.67.

Answer : 2.2) Calculation of Initial After-tax outlay for the machine :

Initial After-Tax Outlay = Cost of New Equipment - After tax salvage Value of old machine

= 800,000 - 325,000

= 475,000

Cost of New Equipment = 800,000

After tax salvage Value of Old Machine = Salvage Value - Tax of Gain on sale

= 340,000 - 15000(Working Note)

= 325,000

Working Note

Gain on Sale = Market Value - Book Value

= 340,000 - 290,000

= 50000

Tax on gain on sale = 50000 * 30%

= 15000